Subject: APO 3.5-m Users Committee minutes 6/14/04

From: Ed Turner

Submitted: Mon, 21 Jun 2004 13:53:29 -0400 (EDT)

Message number: 793 (previous: 792, next: 794 up: Index)

Attendance: Turner, Hawley, Klaene, Bally, Holtzman, Balick, Glazebrook

Absent: Harper, Walterbos, Strauss, Gillespie, York, Green, McMillan

Minutes taken by Jon Holtzman & Ed Turner

Informational item: Suzanne Hawley has been appointed next APO director, will
  be joining meetings this year and taking over 1 January 2005.

NIC-FPS report: Have had some trouble with detector controllers, may be
  that this is a wiring harness or even a chip problem. May require shipping 
  detector to Michigan to isolate the problem. This makes it unlikely that
  NIC-FPS will be delivered to APO at the beginning of August. As is usual
  with issues of this sort, it is difficult to judge the impact on schedule
  until the nature of the problem is securely determined.

Cormass data availability: see apo35-general #792.  If anyone wants to look
  at some of the sample data taken during a test run in early May with
  the U. Va CORMASS low resolution IR spectrograph. Some anecdotal reports 
  suggest that people are very pleased with the data.

U. Va spectrograph: a series of discussions have begun about having the
  group at UVa building a medium resolution IR spectrograph for use at APO.
  Discussions are at a preliminary stage, but U. Va. does seem to have a
  lot of enthusiasm for getting involved at a variety of possible levels;
  these are being discussed.

Summer shutdown: scheduled to start Monday June 28, a bit early because
  of moon phase and some other activities limiting personnel availability
  in August. Current plans are to work on:

   - DIS grating motors, Cryotigers and Cryotiger lines, 
   - DIS slitviewer shutter replacement and dessicant replacement,
   - service UPSs
   - tertiary clamp
   - possibly pull and wash primary mirror
   - vacuum pump instruments 
   - work on power distribution inside computer room
   - replace NA2 guider camera
   - continue work on Apollo project, getting laser powered up and certified
   - some on-sky engineering data, perhaps with Fastcam 
   - preparation for NIC-FPS
   - secondary and tertiary actuators are no longer scheduled to be worked on
     due to manpower issues

NA2 guider: NA2 Photometrics camera will be replaced with an Apogee E6 camera,
   the order has been placed. Work is underway on mounting issues. We hope
   to come out of shutdown with the new guider operational.

We had one recent observing noshow. We may once again consider the possibility
   of having an automatic email notification, perhaps with the option for
   users to exempt themselves from this service.  This will most likely 
   require a small change to the standard ascii-template secheduling 
   request form.

A Futures Committee meeting took place in Cloudcroft at the 10th anniversary
   celebration (which, by the way, was a very nice event!). Don York chairs
   the committee, and almost every institution was represented at the
   first meeting. Discussion was fruitful and ranged over a wide variety
   of topics: from strategies for instrumention in near-term to long-range
   ambitions for ARC. Plans were discussed for a series of meetings, with
   the next meeting in Chicago at the end of summer/beginning of fall.
   All interested people are encouraged to have discussions with the
   representatives, who in fact have been tasked with getting a range
   of information from people at the respective institutions.

Minutes of the 5/10/04 UC meeting were approved without changes.

Next Users Committe meeting date: tentatively August 2nd, but Users 
   committee members are being polled to see if this is the best date.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 793 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to